OpenStudy (anonymous):

f(x)=X(squared)-4x+7 for x>2 g(x)=x-2 for x>2 the function h is such that f=hg and the domain of h is x>0 pleaseee help noww.

OpenStudy (anonymous):

So you are trying to figure out h, right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Ok, h is such that f = hg, which means that if you solve for h, you can figure out what is means in terms of f and g. Given f = hg, what is h? (just in terms of f and g, don't worry about x yet.)

OpenStudy (anonymous):

so it becomes h=f/g? and then i simplify it?

OpenStudy (anonymous):

Yes, h = f/g So, now throw in the expression for f and g.

OpenStudy (anonymous):

So what is h in terms of x?

OpenStudy (anonymous):

f = x^2-4x+7 g = x-2

OpenStudy (anonymous):

yes i divide f by g?? i think that i have to do something with the domain they gave me the teacher said that it is more complicated than that

OpenStudy (anonymous):

Hmmm, I'm not 100% sure. The only problem I see is that f and g are not defined when 0 < x < 2. Gimme a second to think about that. In the meantime, what is f/g when 2 < x?

OpenStudy (anonymous):

if you weree tryin to ask me... i really dont know-.-

OpenStudy (anonymous):

plug in the x's for the expressions of f and g to find h Remember h = f/g

OpenStudy (anonymous):

x as a variable or is there any number that represents it?

OpenStudy (anonymous):

What is f in terms of x? What is g in terms of x? What is h in terms of x? h = f/g

OpenStudy (anonymous):

ahhhh ,okayyyy thanks a lott:))

OpenStudy (anonymous):

So, what is h in terms of x?

OpenStudy (anonymous):

(x-2) squared +5 / x-2

OpenStudy (anonymous):

do i have to simplify it?

OpenStudy (anonymous):

It is simpler if you just leave it in the original form I believe

OpenStudy (anonymous):

yeah i guess so thanks anyway!!!

OpenStudy (anonymous):

No problem. In terms of the domain, I'm still not sure but I think h is just undefined from 0 to 2. But you may want to ask someone else about that. Regardless, you have a major part of the problem completed and understand how to solve similar ones.

OpenStudy (anonymous):

@Hero , do you have a suggestion on this?

hero (hero):

Okay, so do \[\frac{f}{g} = h , x> 2\]

OpenStudy (anonymous):

so what should i substitute x with?

hero (hero):

Which you simply divide \[\frac{x^2 -4x + 7}{x-2}\] You don't need to substitute x with anything yet

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Yes, that's it. But @Hero , is h undefined for 0 < x < 2 ?

OpenStudy (anonymous):

Also, @Hero can you show me how to do the really spiffy math display you used?

hero (hero):

spiffy?

hero (hero):

`\[\frac{x^2 -4x + 7}{x-2}\]`

hero (hero):

Alternatively, just use equation editor

hero (hero):

\[x \ne 2\]

OpenStudy (anonymous):

Wow, that's seems complicated lol. I'd rather just type it out like I am :) haha. Thanks tho!

hero (hero):

It's not complicated

OpenStudy (anonymous):

I understand that x cannot equal 2, but what about x = 1 ? f and g are not defined at x = 1 so therefore h is undefined at x = 1 right?

hero (hero):

To produce \(\huge\frac{a}{b}\) place: `\frac{a}{b}` between `\[\]`

hero (hero):

According to what was originally expressed, h(x) exists whenever x > 0. You shouldn't try to change the constraints of h(x) based on what the constraints of f(x) and g(x) are. Especially when they explicitly tell you what the constraints are. Don't make it more complicated than it should be.

hero (hero):

@teic85

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

However, the student indicated that their teacher said the domains were important and it wasn't as simple as I made it out to be. I had the same opinion as you on the subject. The domains are what they are. To be honest, I feel like the domains are extraneous information to this problem. @Hero

hero (hero):

I agree they are important and recognizing what applies to what is equally important. In that regard, the constraints are clearly specified.

hero (hero):

With h(x), the constraints should be (0,2), (2, ∞)